$SPAD/src/input sint.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool sint.output
)set message test on
)set message auto off
)clear all
--S 1 of 11
min()$SingleInteger
--R
--R
--R   (1)  - 2147483648
--R                                                          Type: SingleInteger
--E 1
--S 2 of 11
max()$SingleInteger
--R
--R
--R   (2)  2147483647
--R                                                          Type: SingleInteger
--E 2
--S 3 of 11
a := 1234 :: SingleInteger
--R
--R
--R   (3)  1234
--R                                                          Type: SingleInteger
--E 3
--S 4 of 11
b := 124$SingleInteger
--R
--R
--R   (4)  124
--R                                                          Type: SingleInteger
--E 4
--S 5 of 11
gcd(a,b)
--R

--R
--R   (5)  2
--R                                                          Type: SingleInteger
--E 5
--S 6 of 11
lcm(a,b)
--R
--R
--R   (6)  76508
--R                                                          Type: SingleInteger
--E 6
--S 7 of 11
mulmod(5,6,13)$SingleInteger
--R
--R
--R   (7)  4
--R                                                          Type: SingleInteger
--E 7
--S 8 of 11
positiveRemainder(37,13)$SingleInteger
--R
--R
--R   (8)  11
--R                                                          Type: SingleInteger
--E 8
--S 9 of 11
And(3,4)$SingleInteger
--R
--R
--R   (9)  0
--R                                                          Type: SingleInteger
--E 9
--S 10 of 11
shift(1,4)$SingleInteger
--R
--R
--R   (10)  16
--R                                                          Type: SingleInteger
--E 10
--S 11 of 11
shift(31,-1)$SingleInteger
--R
--R
--R   (11)  15

--R                                                          Type: SingleInteger
--E 11
)spool
)lisp (bye)

———-

References^{[1] nothing}

6